# Equivariant Machine Learning on Graphs with Nonlinear Spectral Filters

### v1

@@ 1 Introduction
In many fields, such as chemistry [Gilmer et al., 2017], biology [Gaudelet et al., 2021, Barabasi and Oltvai, 2004], social science [Borgatti et al., 2009], and computer graphics [Wang et al., 2019], data can be described by graphs. In recent years, there has been a tremendous interest in the development of machine learning models for graph-structured data [Wu et al., 2020, Bronstein et al., 2017]. This young field, often called graph machine learning (graph-ML) or graph representation learning, has made significant contributions to the applied sciences, e.g., in protein folding, molecular design, and drug discovery [Jumper et al., 2021, Atz et al., 2021, Méndez-Lucio et al., 2021, Stokes et al., 2020], and has impacted the industry with applications in social media, recommendation systems, traffic prediction, computer graphics, and natural language processing, among others.
Geometric Deep Learning (GDL) [Bronstein et al., 2017] is a design philosophy for machine learning models where the model is constructed to inherently respect symmetries present in the data, aiming to reduce model complexity and enhance generalization. By incorporating knowledge of these symmetries into the model, it avoids the need to expend parameters and data to learn them. This inherent respect for symmetries is automatically generalized to test data, thereby improving generalization [Bietti et al., 2021, Petrache and Trivedi, 2024, Elesedy and Zaidi, 2021]. For instance, convolutional neural networks (CNNs) respect the translation symmetries of 2D images, with weight sharing due to these symmetries contributing significantly to their success [Krizhevsky et al., 2012]. Respecting node re-indexing in a scalable manner revolutionized machine learning on graphs [Bronstein et al., 2017, 2021] and has placed GNNs as a main general-purpose tool for processing graph-structured data. Moreover, within GNNs, respecting the 3D Euclidean symmetries of the laws of physics (rotations, reflections, and translations) led to state-of-the-art performance in molecule processing [Duval et al., 2024].
@@ 2 Background
Notation. For $N\in\mathbb{N}$, we denote $[N]=\{1,\ldots,N\}$. We denote matrices by boldface uppercase letter $\mathbf{B}$, vectors (assumed to be columns) by lowercase boldface $\mathbf{b}$, and the entries of matrices and vectors are denoted with the same letter in lower case, e.g., $\mathbf{B}=(b_{i,j})_{i,j\in[N]}$. Let $G=([N],\mathcal{E},\mathbf{A},\mathbf{X})$ be an undirected graph with a node set $[N]$, an edge set $\mathcal{E}\subset[N]\times[N]$, an adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ representing the edge weights, and a node feature matrix $\mathbf{X}\in\mathbb{R}^{N\times d}$ containing $d$-dimensional node attributes. Let $\mathbf{D}$ be the diagonal degree matrix of $G$, where the diagonal element $D_{i,i}$ is the degree of node $i$. Denote by $\bm{\Delta}$ any normal graph shift operator (GSO). For example, $\bm{\Delta}$ could be the combinatorial graph Laplacian or the normalized graph Laplacian given by $\mathbf{L}=\mathbf{D}-\mathbf{A}$ and $\mathbf{N}=\mathbf{D}^{-\frac{1}{2}}\mathbf{L}\mathbf{D}^{-\frac{1}{2}}$, respectively. Let $\mathbf{\Delta}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{\top}$ be the eigendecomposition of $\mathbf{\Delta}$, where $\mathbf{V}$ is the eigenvector matrix and $\mathbf{\Lambda}=\text{diag}(\lambda_{1},\ldots,\lambda_{N})$ is the diagonal matrix with eigenvalues $(\lambda_{i})_{i=1}^{N}$ ordered by $\left\lvert\lambda_{1}\right\rvert\leq\ldots\leq\left\lvert\lambda_{N}\right\rvert$. An eigenspace is the span of all eigenvectors corresponding to the same eigenvalue. Let $\mathbf{P}_{i}=\mathbf{P}_{\bm{\Delta};i}$ denote the projection upon the $i$-th eigenspace of ${\bm{\Delta}}$ in increasing order of $\left\lvert\lambda\right\rvert$. We denote the Euclidean norm by $\left\lVert{\mathbf{X}}\right\rVert_{2}$. We define the channel-wise signal norm $\left\lVert{\mathbf{X}}\right\rVert_{\rm sig}$ of a matrix ${\mathbf{X}}\in{\mathbb{R}}^{N\times d}$ as the vector $\left\lVert{\mathbf{X}}\right\rVert_{\rm sig}=(\left\lVert{\mathbf{X}}_{:,j}\right\rVert_{2})_{j=1}^{d}\in{\mathbb{R}}^{d}$, where ${\mathbf{X}}_{:,j}$ is the $j$-th column on ${\mathbf{X}}$. We abbreviate multilayer perceptrons by MLP.
@@ 2.1 Linear Graph Signal Processing
Spectral GNNs define convolution operators on graphs via the spectral domain. Given a self-adjoint graph shift operator (GSO) $\bm{\Delta}$, e.g., a graph Laplacian, the Fourier modes of the graph are defined to be the eigenvectors $\{\mathbf{v}_{i}\}_{i=1}^{N}$ of $\bm{\Delta}$ and the eigenvalues $\{\lambda_{i}\}_{i=1}^{N}$ are the frequencies. A spectral filter is defined to directly satisfy the “convolution theorem” [Bracewell and Kahn, 1966] for graphs. Namely, given a signal ${\mathbf{X}}\in{\mathbb{R}}^{N\times d}$ and a function $\mathbf{Q}:{\mathbb{R}}\rightarrow{\mathbb{R}}^{d^{\prime}\times d}$, the operator $\mathbf{Q}(\bm{\Delta}):\mathbb{R}^{N\times d}\rightarrow\mathbb{R}^{N\times d^{\prime}}$ defined by
@@
is called a filter. Here, $d^{\prime}$ is the number of output channels. Spectral GNNs, e.g., [Defferrard et al., 2016b, Kipf and Welling, 2016, Levie et al., 2018, Bianchi et al., 2021], are graph convolutional networks where convolutions are via Eq. (1), with a trainable function $\mathbf{Q}$ at each layer, and a nonlinear activation function.
@@ 2.2 Equivariant GNNs
Equivariance describes the ability of functions $f$ to respect symmetries. It is expressed as $f(H_{\kappa}x)=H_{\kappa}f(x)$, where $\mathcal{K}\ni\kappa\mapsto H_{\kappa}$ is an action of a symmetry group $\mathcal{K}$ on the domain of $f$. GNNs [Scarselli et al., 2008, Bruna et al., 2013], including spectral GNNs [Defferrard et al., 2016a, Kipf and Welling, 2016] and subgraph GNNs [Frasca et al., 2022, Bevilacqua et al., 2022], are inherently permutation equivariant w.r.t. the ordering of nodes. This means that the network’s operations are unaffected by the specific arrangement of nodes, a property stemming from passive symmetries [Villar et al., 2023] where transformations are applied to both the graph signals and the graph domain. This permutation equivariance is often compared to the translation equivariance in CNNs [LeCun et al., 1995, 2010], which involves active symmetries [Huang et al., 2024]. The key difference between the two symmetries lies in the domain: while CNNs operate on a fixed domain with signals transforming within it, graphs lack a natural notion of domain translation. To address this, we consider graph functional shifts as the symmetry group, defined by unitary operators that commute with the graph shift operator. This perspective allows for our NLSFs to be interpreted as an extension of active symmetry within the graph context, bridging the gap between the passive and active symmetries inherent to GNNs and CNNs, respectively.
@@ 3 Nonlinear Spectral Graph Filters
In this section, we present new concepts of analysis and synthesis under which the spectral domain is transferrable between graphs. Following these concepts, we introduce new GNNs that are equivariant to functional symmetries – symmetries of the Hilbert space of signals rather than symmetries in the domain of definition of the signal [Lim et al., 2024].
@@ 3.1 Translation Equivariance of CNNs and GNNs
Note that any $\mathbf{T}_{m,n}$ is a unitary operator that commutes with the grid Laplacian $\bm{\Delta}_{R}$, i.e., $\mathbf{T}_{m,n}\bm{\Delta}_{R}=\bm{\Delta}_{R}\mathbf{T}_{m,n}$, and therefore it belongs to the group of all unitary operators $\mathcal{U}_{R}$ that commute with the grid Laplacian $\bm{\Delta}_{R}$. In fact, the space of isotropic convolution operators (with 90o rotation and reflection symmetric filters) can be seen as the space of all normal operators [^1] The operator $\mathbf{B}$ is normal iff $\mathbf{B}^{*}\mathbf{B}=\mathbf{B}\mathbf{B}^{*}$. Equivalently, iff $\mathbf{B}$ has an orthogonal eigendecomposition. that commute with unitary operators from $\mathcal{U}_{R}$ [Cohen and Welling, 2016]. Applying a non-linearity after the convolution retains this equivariance, and hence, we can build multi-layer CNNs that commute with $\mathcal{U}_{R}$. By the universal approximation theorem [Cybenko, 1989, Funahashi, 1989, Leshno et al., 1993], this allows us to approximate any continuous function that commutes with $\mathcal{U}_{R}$.
@@
$\rho\left(\mathbf{Q}(\bm{\Delta})\mathbf{U}\mathbf{X}\right)\neq\mathbf{U}\rho\left(\mathbf{Q}(\bm{\Delta})\mathbf{X}\right).$
This means that multi-layer spectral GNNs do not commute with $\mathcal{U}_{G}$, and are hence not appropriate as approximators of general continuous functions that commute with $\mathcal{U}_{G}$. Instead, we propose in this paper a multi-layer GNN that is fully equivariant to $\mathcal{U}_{G}$, which we show to be universal: it can approximate any continuous graph-signal function (w.r.t. some metric) commuting with $\mathcal{U}_{G}$.
@@ 3.2 Graph Functional Symmetries and Their Relaxations
It is important to note that functional shifts, in general, are not induced from node permutations. Instead, functional shifts are related to the notion of functional maps [Ovsjanikov et al., 2012] used in shape correspondence.
A fundamental challenge with the symmetry group in Def. 3.1 is its lack of transferability between different graphs. Hence, we propose to relax this symmetry group. Let $g_{1},\ldots,g_{S}:\mathbb{R}\rightarrow\mathbb{R}$ be the indicator functions of the intervals $\{[l_{s},l_{s+1}]\}_{s=1}^{S}$, which constitute a partition of the frequency band $[l_{1},l_{S}]\subset\mathbb{R}$. The operators $g_{j}(\bm{\Delta})$, interpreted via functional calculus Eq. (1), are projections of the signal space upon band-limited signals. Namely, $g_{j}(\bm{\Delta})=\sum_{i:\lambda_{i}\in[l_{j},l_{j+1}]}\mathbf{v}_{i}\mathbf{v}_{i}^{\top}$. In our work, we consider filters $g_{j}$ that are supported on the dyadic sub-bands $\left[\lambda_{N}r^{S-j+1},\lambda_{N}r^{S-j}\right]$, where $10<r<$ 1 is the decay rate. See Fig. 3 in App. C for an illustrated example. Note that for $j=1$, the sub-band falls in $[0,\lambda_{N}r^{S-1}]$. The total band $[0,l_{S}]$ is $[0,\lambda_{N}]$.
@@ 3.3 Analysis and Synthesis
We use the terminology of analysis and synthesis, as in signal processing [Mallat, 1989], to describe transformations of signals between their graph and spectral representations. Here, we consider two settings: the eigenspace projections case and the filter bank (of indicators) case.
Spectral Index Case. We first define the analysis and synthesis using the spectral index by
@@
$\mathcal{A}^{\text{ind}}({\bm{\Delta}},{\mathbf{X}})=\Big{(}\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}\Big{)}_{j=1}^{J}\in{\mathbb{R}}^{J\times d}\text{\ \ \ and \ \ \ }\mathcal{S}^{\text{ind}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})=\sum_{j=1}^{J}\mathbf{r}_{j}\ \frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e},$ (2)
@@
respectively, where in synthesis, the product and division are element-wise. Here, $\mathbf{R}=(\mathbf{r}_{j})_{j=1}^{J}\in{\mathbb{R}}^{J\times d}$ are the spectral coefficients to be synthesized, and $0\leq a\leq 1$ and $0<e\ll 1$ are parameters that promote stability. The term index stems from the fact that eigenvalues are treated according to their index when defining the projections $\mathbf{P}_{j}$. Note that the synthesis here differs from classic signal processing as it depends on a given signal on the graph. When treating ${\bm{\Delta}}$ and ${\mathbf{X}}$ as fixed, this synthesis operation is denoted by $\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}^{\text{ind}}(\mathbf{R}):=\mathcal{S}^{\text{ind}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})$. We similarly denote $\mathcal{A}^{\text{ind}}_{{\bm{\Delta}}}({\mathbf{X}}):=\mathcal{A}^{\text{ind}}({\bm{\Delta}},{\mathbf{X}})$.
Filter Bank Case. We also consider and define the analysis and synthesis in the filter bank by
@@
$\mathcal{A}^{\text{val}}({\bm{\Delta}},{\mathbf{X}})=\Big{(}\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}\Big{)}_{j=1}^{K}\in{\mathbb{R}}^{K\times d}\text{\ and \ }\mathcal{S}^{\text{val}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})=\sum_{j=1}^{K}\mathbf{r}_{j}\ \frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e},$ (3)
@@
respectively, where in synthesis, the product and division are element-wise. Here, $\mathbf{R}=(\mathbf{r}_{j})_{j=1}^{K}\in{\mathbb{R}}^{K\times d}$ are the spectral coefficients to be synthesized, and $a,e$ are as before. In App. A.3, we show that the synthesis operator is stably invertible. The term value refers to how eigenvalues are used based on their magnitude when defining the projections $g_{j}({\bm{\Delta}})$. As before, we denote $\mathcal{S}^{\text{val}}_{{\bm{\Delta}},{\mathbf{X}}}$ and $\mathcal{A}^{\text{val}}_{{\bm{\Delta}}}$.

### v2

@@ 1 Introduction
In many fields, such as chemistry [37], biology [36, 3], social science [9], and computer graphics [97], data can be described by graphs. In recent years, there has been a tremendous interest in the development of machine learning models for graph-structured data [98, 13]. This young field, often called graph machine learning (graph-ML) or graph representation learning, has made significant contributions to the applied sciences, e.g., in protein folding, molecular design, and drug discovery [45, 2, 70, 89], and has impacted the industry with applications in social media, recommendation systems, traffic prediction, computer graphics, and natural language processing, among others.
Geometric Deep Learning (GDL) [13] is a design philosophy for machine learning models where the model is constructed to inherently respect symmetries present in the data, aiming to reduce model complexity and enhance generalization. By incorporating knowledge of these symmetries into the model, it avoids the need to expend parameters and data to learn them. This inherent respect for symmetries is automatically generalized to test data, thereby improving generalization [6, 76, 25]. For instance, convolutional neural networks (CNNs) respect the translation symmetries of 2D images, with weight sharing due to these symmetries contributing significantly to their success [54]. Respecting node re-indexing in a scalable manner revolutionized machine learning on graphs [13, 12] and has placed GNNs as a main general-purpose tool for processing graph-structured data. Moreover, within GNNs, respecting the 3D Euclidean symmetries of the laws of physics (rotations, reflections, and translations) led to state-of-the-art performance in molecule processing [24].
@@ 2 Background
Notation. For $N\in\mathbb{N}$, we denote $[N]=\{1,\ldots,N\}$. We denote matrices by boldface uppercase letter $\mathbf{B}$, vectors (assumed to be columns) by lowercase boldface $\mathbf{b}$, and the entries of matrices and vectors are denoted with the same letter in lower case, e.g., $\mathbf{B}=(b_{i,j})_{i,j\in[N]}$. Let $G=([N],\mathcal{E},\mathbf{A},\mathbf{X})$ be an undirected graph with a node set $[N]$, an edge set $\mathcal{E}\subset[N]\times[N]$, an adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ representing the edge weights, and a node feature matrix (also called a signal) $\mathbf{X}\in\mathbb{R}^{N\times d}$ containing $d$-dimensional node attributes. Let $\mathbf{D}$ be the diagonal degree matrix of $G$, where the diagonal element $d_{i,i}$ is the degree of node $i$. Denote by $\bm{\Delta}$ any normal graph shift operator (GSO). For example, $\bm{\Delta}$ could be the combinatorial graph Laplacian or the normalized graph Laplacian given by $\mathbf{L}=\mathbf{D}-\mathbf{A}$ and $\mathbf{N}=\mathbf{D}^{-\frac{1}{2}}\mathbf{L}\mathbf{D}^{-\frac{1}{2}}$, respectively. Let $\mathbf{\Delta}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{\top}$ be the eigendecomposition of $\mathbf{\Delta}$, where $\mathbf{V}$ is the eigenvector matrix and $\mathbf{\Lambda}=\text{diag}(\lambda_{1},\ldots,\lambda_{N})$ is the diagonal matrix with eigenvalues $(\lambda_{i})_{i=1}^{N}$ ordered by $\left\lvert\lambda_{1}\right\rvert\leq\ldots\leq\left\lvert\lambda_{N}\right\rvert$. An eigenspace is the span of all eigenvectors corresponding to the same eigenvalue. Let $\mathbf{P}_{i}=\mathbf{P}_{\bm{\Delta};i}$ denote the projection upon the $i$-th eigenspace of ${\bm{\Delta}}$ in increasing order of $\left\lvert\lambda\right\rvert$. We denote the Euclidean norm by $\left\lVert{\mathbf{X}}\right\rVert_{2}$. We define the channel-wise signal norm $\left\lVert{\mathbf{X}}\right\rVert_{\rm sig}$ of a feature matrix ${\mathbf{X}}\in{\mathbb{R}}^{N\times d}$ as the vector $\left\lVert{\mathbf{X}}\right\rVert_{\rm sig}=(\left\lVert{\mathbf{X}}_{:,j}\right\rVert_{2})_{j=1}^{d}\in{\mathbb{R}}^{d}$, where ${\mathbf{X}}_{:,j}$ is the $j$-th column on ${\mathbf{X}}$ and $0\leq a\leq 1$. We abbreviate multilayer perceptrons by MLP.
@@ 2.1 Linear Graph Signal Processing
Spectral GNNs define convolution operators on graphs via the spectral domain. Given a self-adjoint graph shift operator (GSO) $\bm{\Delta}$, e.g., a graph Laplacian, the Fourier modes of the graph are defined to be the eigenvectors $\{\mathbf{v}_{i}\}_{i=1}^{N}$ of $\bm{\Delta}$ and the eigenvalues $\{\lambda_{i}\}_{i=1}^{N}$ are the frequencies. A spectral filter is defined to directly satisfy the “convolution theorem” [11] for graphs. Namely, given a signal ${\mathbf{X}}\in{\mathbb{R}}^{N\times d}$ and a function $\mathbf{Q}:{\mathbb{R}}\rightarrow{\mathbb{R}}^{d^{\prime}\times d}$, the operator $\mathbf{Q}(\bm{\Delta}):\mathbb{R}^{N\times d}\rightarrow\mathbb{R}^{N\times d^{\prime}}$ defined by
@@
is called a filter. Here, $d^{\prime}$ is the number of output channels. Spectral GNNs, e.g., [21, 49, 61, 5], are graph convolutional networks where convolutions are via Eq. (1), with a trainable function $\mathbf{Q}$ at each layer, and a nonlinear activation function.
@@ 2.2 Equivariant GNNs
Equivariance describes the ability of functions $f$ to respect symmetries. It is expressed as $f(H_{\kappa}x)=H_{\kappa}f(x)$, where $\mathcal{K}\ni\kappa\mapsto H_{\kappa}$ is an action of a symmetry group $\mathcal{K}$ on the domain of $f$. GNNs [83, 14], including spectral GNNs [22, 49] and subgraph GNNs [31, 4], are inherently permutation equivariant w.r.t. the ordering of nodes. This means that the network’s operations are unaffected by the specific arrangement of nodes, a property stemming from passive symmetries [95] where transformations are applied to both the graph signals and the graph domain. This permutation equivariance is often compared to the translation equivariance in CNNs [56, 57], which involves active symmetries [44]. The key difference between the two symmetries lies in the domain: while CNNs operate on a fixed domain with signals transforming within it, graphs lack a natural notion of domain translation. To address this, we consider graph functional shifts as the symmetry group, defined by unitary operators that commute with the graph shift operator. This perspective allows for our NLSFs to be interpreted as an extension of active symmetry within the graph context, bridging the gap between the passive and active symmetries inherent to GNNs and CNNs, respectively.
@@ 3 Nonlinear Spectral Graph Filters
In this section, we present new concepts of analysis and synthesis under which the spectral domain is transferrable between graphs. Following these concepts, we introduce new GNNs that are equivariant to functional symmetries – symmetries of the Hilbert space of signals rather than symmetries in the domain of definition of the signal [64].
@@ 3.1 Translation Equivariance of CNNs and GNNs
Note that any $\mathbf{T}_{m,n}$ is a unitary operator that commutes with the grid Laplacian $\bm{\Delta}_{R}$, i.e., $\mathbf{T}_{m,n}\bm{\Delta}_{R}=\bm{\Delta}_{R}\mathbf{T}_{m,n}$, and therefore it belongs to the group of all unitary operators $\mathcal{U}_{R}$ that commute with the grid Laplacian $\bm{\Delta}_{R}$. In fact, the space of isotropic convolution operators (with 90o rotation and reflection symmetric filters) can be seen as the space of all normal operators [^1] The operator $\mathbf{B}$ is normal iff $\mathbf{B}^{*}\mathbf{B}=\mathbf{B}\mathbf{B}^{*}$. Equivalently, iff $\mathbf{B}$ has an orthogonal eigendecomposition. that commute with unitary operators from $\mathcal{U}_{R}$ [18]. Applying a non-linearity after the convolution retains this equivariance, and hence, we can build multi-layer CNNs that commute with $\mathcal{U}_{R}$. By the universal approximation theorem [19, 34, 58], this allows us to approximate any continuous function that commutes with $\mathcal{U}_{R}$.
@@
$\rho\left(\mathbf{Q}(\bm{\Delta})\mathbf{U}\mathbf{X}\right)\neq\mathbf{U}\rho\left(\mathbf{Q}(\bm{\Delta})\mathbf{X}\right),$
where $\rho$ is any non-linear activation function, e.g., ReLU, Sigmoid, etc.
This means that multi-layer spectral GNNs do not commute with $\mathcal{U}_{G}$, and are hence not appropriate as approximators of general continuous functions that commute with $\mathcal{U}_{G}$ (see App. A for an example illustrating how non-linear activation functions break the functional symmetry). Instead, we propose in this paper a multi-layer GNN that is fully equivariant to $\mathcal{U}_{G}$, which we show to be universal: it can approximate any continuous graph-signal function (w.r.t. some metric) commuting with $\mathcal{U}_{G}$.
@@ 3.2 Graph Functional Symmetries and Their Relaxations
It is important to note that functional shifts, in general, are not induced from node permutations. Instead, functional shifts are related to the notion of functional maps [73] used in shape correspondence and are general unitary operators that are not permutation matrices in general. The value of the functionally translated signal at a given node can be a mixture of the content of the original signal at many different nodes. For example, the functional shift can be a combination of shifts of different frequencies at different speeds. See App. B for illustrations and examples of functional translations.
A fundamental challenge with the symmetry group in Def. 3.1 is its lack of transferability between different graphs. Hence, we propose to relax this symmetry group. Let $g_{1},\ldots,g_{S}:\mathbb{R}\rightarrow\mathbb{R}$ be the indicator functions of the intervals $\{[l_{s},l_{s+1}]\}_{s=1}^{S}$, which constitute a partition of the frequency band $[l_{1},l_{S}]\subset\mathbb{R}$. The operators $g_{j}(\bm{\Delta})$, interpreted via functional calculus Eq. (1), are projections of the signal space upon band-limited signals. Namely, $g_{j}(\bm{\Delta})=\sum_{i:\lambda_{i}\in[l_{j},l_{j+1}]}\mathbf{v}_{i}\mathbf{v}_{i}^{\top}$. In our work, we consider filters $g_{j}$ that are supported on the dyadic sub-bands $\left[\lambda_{N}r^{S-j+1},\lambda_{N}r^{S-j}\right]$, where $10<r<$ 1 is the decay rate. See Fig. 5 in App. F for an illustrated example. Note that for $j=1$, the sub-band falls in $[0,\lambda_{N}r^{S-1}]$. The total band $[0,l_{S}]$ is $[0,\lambda_{N}]$.
@@ 3.3 Analysis and Synthesis
We use the terminology of analysis and synthesis, as in signal processing [68], to describe transformations of signals between their graph and spectral representations. Here, we consider two settings: the eigenspace projections case and the filter bank (of indicators) case. The definition of the frequency domain depends on a given signal ${\mathbf{X}}\in\mathbb{R}^{N\times d}$, where its projections to the eigenspaces of ${\bm{\Delta}}$ are taken as the Fourier modes. Spectral coefficients are modeled as matrices $\mathbf{R}$ that mix the Fourier modes, allowing to synthesize signals of general dimensions.
Spectral Index Case. We first define analysis and synthesis using the spectral index up to frequency $J$. Let $\mathbf{P}_{J+1}=\mathbf{I}-\sum_{j=1}^{J}\mathbf{P}_{j}$ be the orthogonal complement to the first $J$ eigenprojections. Let ${\mathbf{X}}\in\mathbb{R}^{N\times d}$ be the graph signal and $\mathbf{R}\in{\mathbb{R}}^{(J+1)d\times(J+1)\widetilde{d}}$ the spectral coefficients to be synthesized, where $\widetilde{d}$ represents number of output channels. The analysis and synthesis are defined respectively by
@@
$\mathcal{A}^{\text{ind}}({\bm{\Delta}},{\mathbf{X}})=\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J+1}\in{\mathbb{R}}^{(J+1)d}\text{\ and \ }\mathcal{S}^{\text{ind}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})=\left[\frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\right]_{j=1}^{J+1}\mathbf{R},$ (2)
@@
where $0\leq a\leq 1$ and $0<e\ll 1$ are parameters that promote stability, and the power $\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}$ as well as the division in Eq. (2) are element-wise operations on each entry. Here, $\big{[}{\mathbf{P}_{j}{\mathbf{X}}}/{(\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e)}\big{]}_{j=1}^{J+1}\in\mathbb{R}^{N\times(J+1)d}$ denotes concatenation. We remark that the orthogonal complement in the $(J+1)$-th filter alleviates the loss of information due to projecting to the low-frequency bands, and therefore, the full spectral range of the signal can be captured. This is particularly important for heterophilic graphs, which rely on high-frequency components for accurate label representation. The term index stems from the fact that eigenvalues are treated according to their index when defining the projections $\mathbf{P}_{j}$. Note that the synthesis here differs from classic signal processing as it depends on a given signal on the graph. When treating ${\bm{\Delta}}$ and ${\mathbf{X}}$ as fixed, this synthesis operation is denoted by $\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}^{\text{ind}}(\mathbf{R}):=\mathcal{S}^{\text{ind}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})$. We similarly denote $\mathcal{A}^{\text{ind}}_{{\bm{\Delta}}}({\mathbf{X}}):=\mathcal{A}^{\text{ind}}({\bm{\Delta}},{\mathbf{X}})$.
Filter Bank Case. Similarly, we define the analysis and synthesis in the filter bank up to band $g_{K}$ as follows. Let $g_{K+1}({\bm{\Delta}})=\mathbf{I}-\sum_{j=1}^{K}g_{j}(\mathbf{\Delta})$ denote the orthogonal complement to the first $K$ bands. Let ${\mathbf{X}}\in{\mathbb{R}}^{N\times d}$ be the graph signal, and let $\mathbf{R}\in{\mathbb{R}}^{(K+1)d\times(K+1)\widetilde{d}}$ represent the spectral coefficients to be synthesized, where $\widetilde{d}$ refers to the general dimension. The analysis and synthesis in the filter bank case are defined by
@@
$\mathcal{A}^{\text{val}}({\bm{\Delta}},{\mathbf{X}})=\left(\left\lVert g_{i}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{K+1}\in{\mathbb{R}}^{(K+1)d}\text{ and }\mathcal{S}^{\text{val}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})=\left[\frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\right]_{j=1}^{K+1}\mathbf{R},$ (3)
@@
respectively, where $a,e$ are as before. Here, $\big{[}{g_{j}({\bm{\Delta}}){\mathbf{X}}}/{(\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e)}\big{]}_{j=1}^{K+1}\in\mathbb{R}^{N\times(K+1)d}$. The term value refers to how eigenvalues are used based on their magnitude when defining the projections $g_{j}({\bm{\Delta}})$. As before, we denote $\mathcal{S}^{\text{val}}_{{\bm{\Delta}},{\mathbf{X}}}$ and $\mathcal{A}^{\text{val}}_{{\bm{\Delta}}}$.
In App. C.1, we present a special case of diagonal synthesis where $\widetilde{d}=d$. In App. D.3, we show that the diagonal synthesis is stably invertible.